OpenStudy (anonymous):

a hydrogen atom has adiameter of about 10nm a) express its diameter in millimeters b) express this diameter in micrometers

OpenStudy (anonymous):

I see that there are 1 million nanometers in one millimeter, and there are one thousand nanometers in one micrometer. Use these conversion factors to solve.

OpenStudy (aum):

1 mm = \(\large 10^{-3}\) m 1 nm = \(\large 10^{-9}\) m divide: 1 mm / 1 nm = \(\large 10^{-3} / 10^{-9} = 10^{-3+9} = 10^6\) 1 mm = \(\large 10^6 \) nm or 1 nm = \(\large 10^{-6} \) mm